OpenStudy (girlgoyleh.):

whats the coefficient of the term -3gh?

OpenStudy (mathmate):

the coefficient of a term includes the sign and any constant (numbers, \(\pi\), etc., i.e. the term stripped of the variables. g and h are rarely used as variables, but they look like variables, so if you remove the variables, what's left will be the coefficient. For example, -3 is the coefficient of -3xyz. \(-\sqrt{48}\pi \) is the coefficient of \(-\sqrt{48}\pi x^2y\)

OpenStudy (girlgoyleh.):

So, always the coefficient is the number?

OpenStudy (mdilley):

the coefficient is -3, because gh are variables.